# SSF-Net: Spatial-Spectral Fusion Network with Spectral Angle Awareness for Hyperspectral Object Tracking

### v1

@@ III-B Bi-stream Feature Extraction Backbone
Figure 3: Illustration of our proposed $S^{2}$. It consists of several $S^{2}$ with the residual connection.
@@
To extract the robust features, a bi-stream feature extraction backbone is proposed, which consists of the HS stream and the RGB stream. Most existing HS tracking methods utilized the off-the-shelf models in the RGB domain, such as VGG, ResNet, etc, to extract the HS features. However, HS images constitute a data cube enriched with information across numerous spectral bands, which can offer a wealth of multidimensional information, and the backbone network pre-trained from three-channel RGB images is insufficient to extract robust spatial-spectral features.
@@
Thus, we propose a feature extraction backbone $S^{2}$ for HS-stream. Its main structure is shown in the Figure 3. The size of input images is $H\times W\times B$, where $H$ and $W$ denote the height and width of the input images, and $B$ denotes the number of spectral bands. The $S^{2}$ consists of four feature extraction stages. Each stage captures different feature responses to the input data. In the early stages, the network focuses on capturing details and local features, while in the later stages, as the network deepens, it captures higher-level semantic information. So, $S^{2}$ can extract discriminative features across multiple layers, ranging from shallow to deep. The first two stages, which are referred to as “Stage 1” and “Stage 2”, include $S^{2}$ and downsampling layers. The purpose of these stages is to ensure that the feature size is compatible with the RGB stream for feature fusion and to capture texture information. The output resolutions of these stages are $\frac{H}{2}\times\frac{W}{2}\times C$ and $\frac{H}{4}\times\frac{W}{4}\times 2C$, respectively. $C$ denotes a channel dimension, which equals 256 in the SSF-Net. The subsequent stages, Stage 3 and Stage 4, repeat the process without downsampling and produce output with the resolutions of $\frac{H}{4}\times\frac{W}{4}\times 4C$ and $\frac{H}{4}\times\frac{H}{4}\times 8C$, respectively. Each $S^{2}$ within the stages consists of two $S^{2}$ layers, along with batch normalization (BN) and ReLU activation layers. As shown in Figure 3, the $S^{2}$ layer utilizes 2-D depth-wise separable convolution to extract spatial features and 3-D convolution to extract multi-spectrum features. The sliding 3-D convolutional kernels establish contextual relationships between multiple continuous spectral bands and capture spectral signatures of substances. This allows the $S^{2}$ to generate a joint spatial-spectral hierarchical representation. Additionally, the use of skip connections in the $S^{2}$ helps mitigate issues such as vanishing gradients and facilitates information flow across different layers. In the RGB branch of our network, we utilize ResNet-50 as the feature extraction backbone. This backbone is pre-trained on RGB tracking datasets, specifically those used in the Siamese tracking framework proposed by Chen et al. [21]. By leveraging the pre-trained knowledge in the RGB domain, we can transfer this knowledge to the task of HS and RGB fusion tracking. This allows us to benefit from the expertise and insights gained from previous research in RGB tracking, and enhance the performance and effectiveness of our proposed network.
@@ III-C Spectral Attention Fusion Module
In recent HS object tracking methods, feature fusion has been achieved through a spectral attention mechanism. However, most of the feature refinement is carried out within modalities, without fully considering the cross-modality interaction relationships and the complementary nature of modality features. To address these limitations, SAFM is proposed for cross-modality feature fusion, while modeling the feature relationships within and between modalities. As shown in Figure 4, a global average pooling is first used to aggregate the spatial information into a channel-wise global feature, which is useful for capturing the overall presence of features in the input features. Assuming that $R_{f}\in{\mathbb{R}^{H\times W\times C}}$ and $H_{f}\in{{\mathbb{R}}^{H\times W\times C}}$ are the extracted features from RGB and HS images, respectively, the global average pooling is utilized to generate features $R_{a}\in{\mathbb{R}^{1\times C}}$ and $H_{a}\in{\mathbb{R}^{1\times C}}$. Then, intra-modality spectral embedding and inter-modality spectral embedding are achieved through 1-D convolution with different parameters. To capture local cross-channel interaction information and account for the similarity of adjacent spectra, we fix the size of the 1-D convolutional kernel. This allows us to effectively model the relationships between channels. It is formulated by
@@
where $C1D_{1}$ and $C1D_{2}$ denote two different $1\times 1$ convolution layers, respectively. For the intra-modality attention, with the spectral embedding features, the spectral attention weights are obtained through the $Sigmoid$ activation function, which is formulated by
@@
The intra-modality refined features can be obtained from the product of input features and attention weights, which are formulated by:
@@
Then the inter-modality refined features between the RGB and HS modalities are obtained by the following formula:
@@ III-D Spectral Angle Awareness Module
This section introduces the proposed spectral angle awareness module, including the spectral angle awareness head, classification head, and localization head. The existing methods only use the classification and localization branches in the existing RGB domain, and calculate the depth-wise similarity of each feature [36] to predict the position and size of the object. They do not fully utilize the spectral information in HS and lack modeling of spectral relationships, which will affect the robustness of the features. To solve this problem, a spectral angle awareness module is proposed.
1) Spectral angle mapping revisit. The spectral angle mapping (SAM), initially introduced by [15], is a widely used measure in remote sensing [37, 38] to assess the similarity between spectral feature vectors. In the HS object detection field, SAM calculates the similarity between an object pixel and a test pixel by first calculating their corresponding mapping matrix. This matrix is then used to compute the cosine of the angle between the spectral vectors of the pixels. The cosine value serves as an effective measure of similarity, with a higher cosine indicating a closer match to the spectral signature of the objects being compared. The specific details are as follows.
Assuming that the spectral feature $\hat{S}\in{{\mathbb{R}}^{H\times W\times C}}$ extracted by machine learning algorithms is a three-dimensional data cube, which is first transformed into a two-dimensional matrix $S\in{{\mathbb{R}}^{HW\times C}}$. The original mapping matrix $M$ is obtained by:
@@
${M}^{-1/2}=Q{V^{-1/2}}Q,$ (7)
where $Q$ is an orthogonal identity matrix, and $V$ is the eigenvalue diagonal matrix of $M$. With the mapping matrix ${M}^{-1/2}$, the spectrum $t$ of an object and sparse tensor $S$ are projected into the same subspace. The cosine similarity is formulated as:
@@
where $(i,j)$ represents the pixel location of feature maps $S$ to be detected. $\cos_{(i,j)}^{\wedge}$ is the spectral angle cosine similarity between the spectrum of the object $t$ and feature map $S$. With the cosine similarity, the influence of background clutter is released and the performance of object detection can be improved.
Figure 5: The architecture of the SAAM includes the CLS branch, LOC branch, and SAA branch.
2) Spectral angle awareness branch. Inspired by the spectral angle calculation used to assess depth similarity between pixels, the spectral angle awareness module is proposed. The structure of SAAM is shown in Figure 5. To improve upon the existing formulation, we have reformulated Equation (8) and presented a simplified formula for computing similarity in the SAM. This simplified formula allows for a more efficient and effective calculation, which is formulated as follows:
@@
where $z$ and $x$ represent TI and SI of the HS modality. $f(z)$, $f(x)$ are the feature maps extracted by the same $f$, which is a series of feature extraction and regularization functions. Here, $f$ represents the feature extraction network $S^{2}$. Indeed, the cosine similarity is calculated by taking the dot product of the mapped HS feature vectors and normalizing it by the magnitude of each vector. In the field of HS tracking, both the classification and localization heads rely on region-level similarity assessments to make their predictions. The features of the TI and SI extracted by the tracker are denoted as $f(z)$ and $f(x)$, respectively. To perform similarity assessments, the feature map $f(x)$ undergoes depth-wise separable convolutions using $f(z)$ as the convolution kernel. This operation is represented by the following formula:
@@
where *** represents depth-wise cross-correlation layer [36]. It can be seen as an extended region-level version of the SAM, all of which aim to obtain similarity through vector operations. The idea of spectral angle can seamlessly integrate into the tracking prediction head. To achieve this integration, an embedding layer is employed to project the feature maps $f(z)$ and $f(x)$ onto a subspace. This subspace projection enables the calculation of cosine similarity between the two feature maps. It is formulated by:
@@
where $E_{z}$ and $E_{x}$ denote the embedding layers used by $f(z)$ and $f(x)$, respectively. The embedded feature $f_{z}^{embed}$ is utilized as a sliding convolution kernel to convolve with $f_{x}^{embed}$, thereby generating the output map. Consequently, this procedure facilitates the calculation of the region-level affinity matrix between $f_{z}^{embed}$ and $f_{x}^{embed}$. Subsequently, a predicted map of spectral angular affinity $M_{SAA}$ is derived using a $1\times 1$ convolutional layer. The detailed formula is presented as:
@@
To train the SAAM more effectively, the SAAL is proposed. The positive and negative sample regions as [21] are constructed to transfer the triplet optimization [39] to the HS tracking. The core idea behind triplet loss is to take three data samples at a time, typically referred to as an “anchor”, a “positive”, and a “negative”. The anchor and positive samples belong to the same class or depict the same object, while the negative samples come from a different class. Therefore, the SAAL is designed to maximize the similarity between the anchor and the positive while minimizing the similarity between the anchor and the negative. As shown in Figure 6, the ground-truth bounding box of SI is marked with the red box, we construct two ellipses $E1$ and $E2$. The position $(i,j)$ inside $E1$ is marked as positive $p$, and the position $(i,j)$ outside $E2$ is marked as negative $n$. We use the features of TI extracted by the tracker as anchors to form positive pairs with positive regions in SI and negative pairs with negative regions in SI. The SAAL is formulated by:
@@
where $p\in E1$ denotes the point inside $E1$ of the $M_{SAA}$, and ${n\in\complement(E2)}$ denote the complementary point of $E2$, i.e., outside $E2$. $a$ denotes the template image. $Sim(a,p)$ and $Sim(a,n)$ denote the cosine similarities of positive pairs and negative pairs, respectively. During the optimization process, the network learns to increase the similarity between positive sample pairs, while decreasing the similarity between negative sample pairs. This means that the angle between the feature vectors of positive sample pairs in the feature space becomes smaller, while the angle between the feature vectors of negative sample pairs becomes larger. This intuitive understanding is illustrated in Figure 1(c). During inference, SAAM assists the trackers in enhancing the response of objects within TI as it appears in SI. This enhancement facilitates the precise localization of the object within the region of similarity shared between TI and SI.
@@
where $\alpha$, $\beta$, and $\gamma$ denote the trade-off parameters. Here, we empirically set $\alpha=1$, $\beta=2$, and $\gamma=1$ for all experiments. Then, to ensemble the prediction from HS and RGB modalities, a weighted prediction method is adopted to learn the weights of the two modalities in a self-adaptive way. The weighted fusion prediction $M_{CLS}$ and $M_{LOC}$ are obtained from:
@@
where ${\lambda_{1}}$ and ${\lambda_{2}}$ denote the trainable parameters, respectively. With the weighted prediction, the CLS and LOC maps integrate the information from two branches, thereby enhancing the accuracy of object localization. Finally, considering the widespread utilization of ensemble prediction in existing methods, we aggregate the prediction maps from the final three stages of both the RGB and HS backbone networks. Furthermore, through the joint calculation of the loss function, the proposed method allows for the simultaneous prediction of the class and location of objects during the inference stage. This integrated approach enhances the efficiency and effectiveness of object prediction, providing more accurate and reliable results.
@@ IV-A Experiment Settings
Datasets. The HOTC dataset [6] is used to validate the effectiveness of the proposed method. The dataset is divided into a training set comprising 40 videos and a testing set consisting of 35 videos. To capture different objects, both RGB and HS cameras are employed for simultaneous recording, resulting in each video containing three types of modality data: RGB, HS, and a generated false-color image. The HS image cubes within the dataset consist of 16 spectral bands. Furthermore, the dataset contains 11 challenging factors, including background clutter (BC), deformation (DEF), fast motion (FM), illumination variation (IV), in-plane rotation (IPR), low resolution (LR), motion blur (MB), out-of-occlusion (OCC), plane rotation (OPR), out-of-view (OV), and scale variation (SV).
@@
Implementation Details. All experiments are conducted using an NVIDIA RTX 4090 GPU. For the RGB branches, we utilize a pre-trained ResNet-50 network sourced from training on the COCO [41], ImageNet VID [42], and YouYube-BB [43] dataset as the backbone network. The SSF-Net is not initialized with any pre-trained parameters. During the training phase, the parameters of the RGB stream are kept fixed, while only the parameters of the HS stream are subject to training. We set the batch size to 30 and the learning rate is initialized as 0.005, applying a warm-up learning rate adjustment strategy. The weight decay and momentum are configured at 0.0001 and 0.9, respectively.
IV-B Comparison with State-of-the-arts
In this subsection, we conduct multiple experiments to demonstrate the effectiveness of our method, including performance comparisons with advanced HS trackers, performance comparisons with advanced RGB trackers, attribute-based comparisons, and visualization analysis.
Comparison with HS Trackers. We first compare the proposed SSF-Net with advanced HS trackers, including band regrouping-based methods SEE-Net, BAE-Net, SST-Net, and band aware-based method SPIRIT. BAE-net initially learned spectral weights in the form of channel attention, and it regrouped the HS cubes to multiple three-channel false-color images based on the spectral weight, thereby transferring existing RGB trackers to process HS tracking. SEE-Net learned the weights between bands in the form of an autoencoder. SST-Net built a temporal-aware false-color image generation framework. SPIRIT utilized both channel attention and an autoencoder-like structure for band awareness, while also considering both intra-band and inter-band interactions. It is worth noting that we reproduced SEE-Net and the results of other trackers can be obtained from the corresponding paper.
TABLE I: overall performance comparison of HS and visual tracker. The HS trackers deal with the HSIs and the visual tracker runs on the false-color images. We use the red and blue fonts to highlight the top two values.
@@ IV-B Comparison with State-of-the-arts
Figure 7: The visualization of heatmaps on the HOTC dataset. The red color represents high response while the blue color represents low response.

### v2

@@ III-B Bi-stream Feature Extraction Backbone
Figure 3: Illustration of our proposed $S^{2}$. It consists of several $S^{2}$ with the residual connection. Each $S^{2}$ contains $S^{2}$, a Batch Normalization layer, and ReLU layers, where $S^{2}$ consists of parallel spatial feature extraction branches and spectral feature extraction branches.
@@
To extract the robust features, a bi-stream feature extraction backbone is proposed, which consists of the HS stream and the RGB stream. Most existing HS tracking methods utilize the off-the-shelf models in the RGB domain, such as VGG, ResNet, etc, to extract the HS features. However, an HS image is a data cube enriched with information across numerous spectral bands, which can offer a wealth of multi-dimensional information, and the backbone network pre-trained from three-channel RGB images is insufficient to extract robust spatial-spectral features.
@@
Thus, we propose a feature extraction backbone $S^{2}$ for HS-stream. Its main structure is shown in Figure 3. The size of input images is $H\times W\times B$, where $H$ and $W$ denote the height and width of the input images, respectively, and $B$ denotes the number of spectral bands. The $S^{2}$ consists of four feature extraction stages. Each stage captures different feature responses to the input data. In the early stages, the network focuses on capturing details and local features, while in the later stages, as the network is deepened, it captures higher-level semantic information. Thus, $S^{2}$ can extract discriminative features across multiple layers, ranging from shallow to deep. The first two stages, which are referred to as “Stage 1” and “Stage 2”, include $S^{2}$ and downsampling layers. The objective of these stages is to ensure that the feature size is compatible with the RGB stream for feature fusion and to capture texture information. The output resolutions of these stages are $\frac{H}{2}\times\frac{W}{2}\times C$ and $\frac{H}{4}\times\frac{W}{4}\times 2C$, respectively. Here, $C$ denotes a channel dimension, which equals 256 in the SSF-Net. The subsequent stages, Stage 3 and Stage 4, repeat the process without downsampling and produce output with the resolutions of $\frac{H}{4}\times\frac{W}{4}\times 4C$ and $\frac{H}{4}\times\frac{H}{4}\times 8C$, respectively. Each $S^{2}$ within the stages consists of two $S^{2}$ layers, along with batch normalization (BN) and ReLU activation layers. As shown in Figure 3, the $S^{2}$ layer utilizes 2-D depth-wise separable convolution to extract spatial features and 3-D convolution to extract multi-spectrum features. The sliding 3-D convolutional kernels establish contextual relationships between multiple continuous spectral bands and capture spectral signatures of substances. This allows the $S^{2}$ to generate a joint spatial-spectral hierarchical representation. Additionally, the use of skip connections in the $S^{2}$ helps mitigate issues such as vanishing gradients and facilitates information flow across different layers. In the RGB branch of our network, we utilize ResNet-50 as the feature extraction backbone. This backbone is pre-trained on RGB tracking datasets, specifically those used in the Siamese tracking framework proposed by Chen et al. [27]. By leveraging the pre-trained knowledge in the RGB domain, we can transfer this knowledge to the task of HS and RGB fusion tracking. The benefit from the expertise and insights gained from previous research in RGB tracking can enhance the performance and effectiveness of the proposed network.
@@ III-C Spectral Attention Fusion Module
In recent HS object tracking methods, feature fusion has been achieved through a spectral attention mechanism. However, most of the feature refinement is carried out within modalities, without fully considering the cross-modality interaction relationships and the complementary nature of modality features. To address these limitations, SAFM is designed to model the feature relationships within and between modalities. As shown in Figure 4, a global average pooling is first used to aggregate the spatial information into a channel-wise global feature, which is useful for capturing the overall presence of features in the input features. Let $R_{f}\in{\mathbb{R}^{H\times W\times C}}$ and $H_{f}\in{{\mathbb{R}}^{H\times W\times C}}$ be the extracted features from RGB and HS images, respectively and the global average pooling is utilized to generate features $R_{a}\in{\mathbb{R}^{1\times C}}$ and $H_{a}\in{\mathbb{R}^{1\times C}}$. Then, intra-modality spectral embedding and inter-modality spectral embedding are achieved through 1-D convolution with different parameters. To capture local cross-channel interaction information and account for the similarity of adjacent spectra, the size of the 1-D convolutional kernel is fixed. This allows effective modeling of the relationships between channels. It is formulated by
@@
where $C1D_{1}$ and $C1D_{2}$ are two different $1\times 1$ convolution layers. For the intra-modality attention, with the spectral embedding features, the spectral attention weights are obtained through the $Sigmoid$ activation function, which is formulated by
@@
The intra-modality refined features can be obtained from the product of input features and attention weights, which are expressed as:
@@
Then, the inter-modality refined features between the RGB and HS modalities are obtained by the following formula:
@@ III-D Spectral Angle Awareness Module
The existing methods only use the classification and localization branches in the existing RGB domain and calculate the depth-wise similarity of each feature [44] to predict the position and size of the object. To fully utilize the spectral information in HS, a spectral angle awareness module is proposed.
1) Spectral angle mapping revisit. The spectral angle mapping (SAM), initially introduced by [17], is a widely used measure in remote sensing [45, 46] to assess the similarity between spectral feature vectors. In the HS object detection field, SAM calculates the similarity between an object pixel and a test pixel by first calculating their corresponding mapping matrix. This matrix is then used to compute the cosine of the angle between the spectral vectors of the pixels. The cosine value serves as an effective measure of similarity, with a higher cosine value indicating a closer match to the spectral signature of the objects being compared. The specific details are as follows.
Assuming that the spectral feature $\hat{S}\in{{\mathbb{R}}^{H\times W\times C}}$ is a three-dimensional data cube, and it is first transformed into a two-dimensional matrix $S\in{{\mathbb{R}}^{HW\times C}}$. The original mapping matrix $M$ is obtained by:
@@
${M}^{-1/2}=Q{V^{-1/2}}Q^{T},$ (7)
where $Q$ is a unitary matrix, and $V$ is the eigenvalue diagonal matrix of $M$. With the mapping matrix ${M}^{-1/2}$, the spectrum $t$ of an object and a sparse tensor $S$ are projected into the same subspace. The cosine similarity between the spectrum of the object $t$ and feature map $S$ is formulated as:
@@
where $(i,j)$ represents the pixel location of feature maps $S$ to be detected. With the cosine similarity, the influence of background clutter can be mitigated, and the performance of object detection can be improved.
Figure 5: The architecture of the SAAM includes the CLS branch, LOC branch, and SAA branch. The CLS and SAA branches predict the position of the objects together, and the LOC branch predicts the size of the objects.
2) Spectral angle awareness branch. The structure of SAAM is shown in Figure 5. To improve upon the existing formulation, we have reformulated Equation (8) and presented a simplified formula for computing similarity in the SAM. This simplified equation allows for a more efficient and effective calculation, which is given by:
@@
where $z$ and $x$ represent TI and SI of the HS modality, respectively, $f(z)$ and $f(x)$ are the feature maps extracted by the same $f$, which is a series of feature extraction and regularization functions. In this work, $f$ is the feature extraction network $S^{2}$. In the field of HS tracking, both the classification and localization heads rely on region-level similarity assessments to make their predictions. The features of the TI and SI extracted by the tracker are denoted as $f(z)$ and $f(x)$, respectively. To perform similarity assessments, the feature map $f(x)$ undergoes depth-wise separable convolutions using $f(z)$ as the convolution kernel. This operation is given by:
@@
where * represents depth-wise cross-correlation layer [44]. It can be seen as an extended region-level version of the SAM, all of which aim to obtain similarity through vector operations. The idea of spectral angle can seamlessly integrate into the tracking prediction head. To achieve this integration, an embedding layer is employed to project the feature maps $f(z)$ and $f(x)$ onto a subspace. This subspace projection enables the calculation of cosine similarity between the two feature maps, expressed as:
@@
where $E_{z}$ and $E_{x}$ denote the embedding layers used by $f(z)$ and $f(x)$, respectively. The embedded feature $f_{z}^{embed}$ is utilized as a sliding convolution kernel to convolve with $f_{x}^{embed}$, resulting in an output map. Consequently, this procedure facilitates the calculation of the region-level affinity matrix between $f_{z}^{embed}$ and $f_{x}^{embed}$. Subsequently, a predicted map of spectral angular affinity $M_{SAA}$ is derived using a $1\times 1$ convolutional layer. The detailed formula is presented as:
@@
To train the SAAM more effectively, the SAAL is proposed. The positive and negative sample regions, as [27], are constructed to transfer the triplet optimization [47] to the HS tracking. The core idea behind triplet loss is to take three data samples at a time, typically referred to as an “anchor”, a “positive”, and a “negative”. The anchor and positive samples belong to the same class or depict the same object, while the negative samples come from a different class. Therefore, the SAAL is designed to maximize the similarity between the anchor and the positive while minimizing the similarity between the anchor and the negative. As shown in Figure 6, the ground-truth bounding box of SI is marked with the red box, and two ellipses $E1$ and $E2$ are constructed. The position $(i,j)$ inside $E1$ is marked as positive $p$, and the position $(i,j)$ outside $E2$ is marked as negative $n$. We use the features of TI extracted by the tracker as anchors to form positive pairs with positive regions in SI and negative pairs with negative regions in SI. The SAAL is formulated by:
@@
where $p\in E1$ denotes the point inside $E1$ of the $M_{SAA}$, and ${n\in\complement(E2)}$ is the complementary point of $E2$, i.e., outside $E2$. $a$ denotes the template image, and $Sim(a,p)$ and $Sim(a,n)$ are the cosine similarities of positive pairs and negative pairs, respectively. During the optimization process, the network learns to increase the similarity between positive sample pairs while decreasing the similarity between negative sample pairs. This means that the angle between the feature vectors of positive sample pairs in the feature space becomes smaller, while the angle between the feature vectors of negative sample pairs becomes larger. This intuitive understanding is illustrated in Figure 1(c). During inference, SAAM assists the trackers in enhancing the response of objects within TI as it appears in SI. This enhancement facilitates the precise localization of the object within the region of similarity shared between TI and SI.
@@
where $\alpha$, $\beta$, and $\gamma$ denote the trade-off parameters. Here, we set $\alpha=1$, $\beta=2$ following [14], and $\gamma=1$ is set to 1 for all experiments. Then, to ensemble the prediction from HS and RGB modalities, a weighted prediction method is adopted to learn the weights of the two modalities in a self-adaptive way. The weighted fusion prediction $M_{CLS}$ and $M_{LOC}$ are obtained from:
@@
where ${\lambda_{1}}$ and ${\lambda_{2}}$ are the trainable parameters. With the weighted prediction, the CLS and LOC maps integrate the information from two branches to enhance the accuracy of object localization. Finally, considering the widespread utilization of ensemble prediction in existing methods, we aggregate the prediction maps from the final three stages of both the RGB and HS backbone networks. Furthermore, through the joint calculation of the loss function, the proposed method can simultaneously predict the class and location of objects during the inference stage. This integrated approach enhances the efficiency and effectiveness of object prediction.
@@ IV-A Experiment Settings
Datasets. The HOTC-2020 dataset [8] is used to validate the effectiveness of the proposed method. The dataset is divided into a training set comprising 40 videos and a testing set consisting of 35 videos. To capture different objects, both RGB and HS cameras are employed for simultaneous recording, resulting in each video containing three types of modality data: RGB, HS, and generated false-color images. The HS image cubes within the dataset consist of 16 spectral bands. Furthermore, the dataset contains 11 challenging factors, including background clutter (BC), deformation (DEF), fast motion (FM), illumination variation (IV), in-plane rotation (IPR), low resolution (LR), motion blur (MB), out-of-occlusion (OCC), plane rotation (OPR), out-of-view (OV), and scale variation (SV).
The HOTC-2024 is developed for the 2024 Hyperspectral Object Tracking Challenge and features three distinct band configurations: 15 bands (RedNIR), 16 bands (VIS), and 25 bands (NIR). Each video frame is stored as a 2D image and can be converted into a 3D image cube using publicly available algorithms. The dataset includes 218 hyperspectral videos for training and 118 for validation, with each frame annotated with a bounding box for the tracked object. We conduct experiments using datasets of 15 bands, labeled as HOTC-2024-RedNIR.
The BihoT [18] is the first dataset used for hyperspectral camouflage object tracking, with 41,912 images captured by an IMEC XIMEA MQ022HG-IM-SM5X5-NIR camera in the 25 bands, including 22,992 images in the training set and 18,920 images in the testing set. The dataset is annotated with 9 attributes, including Background Clutter (BC), Fast Motion (FM), In-Plane Rotation (IPR), Illumination Variation (IV), Low Resolution (LR), Occlusion (OCC), Out-of-Plane Rotation (OPR), Spectral Consistency (SC), and Spectral Variation (SV).
@@
Implementation Details. All experiments are conducted using an NVIDIA RTX 4090 GPU. For the RGB branches, we utilize a pre-trained ResNet-50 network sourced from training on the COCO [49], ImageNet VID [50], and YouTube-BB [51] dataset as the backbone network. The SSF-Net is not initialized with any pre-trained parameters. During the training phase, the parameters of the RGB stream are fixed, while only the parameters of the HS stream are subject to training. We set the batch size to 30, and the learning rate is initialized as 0.005, applying a warm-up learning rate adjustment strategy. The weight decay and momentum are configured at 0.0001 and 0.9, respectively.
TABLE I: Overall performance comparison of HS and visual tracker. The HS trackers deal with the HSIs, and the visual tracker runs on the false-color images. We use the red and blue fonts to highlight the top two values.
@@
TABLE II: Overall performance comparison of HS trackers on the HOTC-2024-RedNIR and BihoT datasets. We use the red and blue fonts to highlight the top two values.
Datasets Metrics SEE-Net[13] SiamHYPER[14] SENSE[55] MMF-Net[56] SSF-Net
HOTC-2024-RedNIR AUC 0.332 0.328 0.351 0.324 0.357
DP_20 0.433 0.428 0.469 0.424 0.463
FLOPs 297.99G 121.70G 287.74G 656.67G 83.37G
FPS 12.06 12.62 11.55 3.75 13.05
BihoT AUC 0.315 0.403 0.454 0.475 0.497
DP_20 0.422 0.539 0.576 0.580 0.668
@@
Figure 7: The visualization of heatmaps on the HOTC-2020 dataset. The red color represents high response while the blue color represents low response.
IV-B Comparison with State-of-the-arts
Comparison with HS Trackers. We first compare the proposed SSF-Net with advanced HS trackers, including band regrouping-based methods SEE-Net, BAE-Net, SST-Net, and band aware-based method SPIRIT. BAE-net initially learned spectral weights in the form of channel attention, and it regrouped the HS cubes to multiple three-channel false-color images based on the spectral weight, thereby transferring existing RGB trackers to process HS tracking. SEE-Net learned the weights between bands in the form of an autoencoder. SST-Net built a temporal-aware false-color image generation framework. SPIRIT utilized both channel attention and an autoencoder-like structure for band awareness while also considering both intra-band and inter-band interactions. It is worth noting that we reproduced SEE-Net, and the results of SPIRIT are not available, thus we obtain the accuracy from the corresponding paper directly without attribute-based and visualized tracking results.